OpenStudy (anonymous):

5^3= ?? Google is confusing me

OpenStudy (mimi_x3):

125

OpenStudy (anonymous):

125

OpenStudy (saifoo.khan):

5 * 5 * 5 = 125

OpenStudy (anonymous):

5*5*5 = 125

OpenStudy (anonymous):

thxx guys

OpenStudy (anonymous):

5*5*5=125

OpenStudy (anonymous):

well how is google confusing you ?

OpenStudy (anonymous):

type "5^3 =" on google and you will get the answer 5^3 = 125 More about calculator.